OpenStudy (anonymous):

derivative of cos(x^2+3)-x^3

OpenStudy (agreene):

-3x^2-2xsin(x^2+3)

OpenStudy (agreene):

Chain rule.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Don't forget to give him a medal ;)

OpenStudy (lalaly):

\[(-\sin(x^2+3) \times 2x )- 3x^2= -2xsin(x^2+3) -3x^2\]